OpenStudy (anonymous):

Which of the following systems of equations has infinitely many solutions? A. 3x − 3y = 11; 6y = 9x − 22 B. 5y = -9x − 6; -27x + 15y = 15 C. 4x − 7y = 7; -14y = 8x − 14 D. -4x + 2 = 3y; 12y − 8 = -16x

OpenStudy (anonymous):

@gypsy1274 help please

OpenStudy (anonymous):

@dpasingh

OpenStudy (anonymous):

C. 4x − 7y = 7; -14y = 8x − 14

OpenStudy (anonymous):

write in the form a1x+b1y=c1 and a2x+b2y=c2 \[if \frac{ a1 }{ a2 }=\frac{ b1 }{ b2 }=\frac{ c1 }{ c2 }\] then there are infinite solutions